# Scientific Large Language Models: A Survey on Biological & Chemical Domains

### v1

@@ 5.2. Datasets
  • TAPE (Rao et al., [n. d.]). The Tasks Assessing Protein Embeddings (TAPE) dataset is a collection of five biologically relevant semi-supervised learning tasks spanning different domains of protein biology, including secondary structure prediction, contact prediction, remote homology prediction, stability prediction, and fluorescence prediction. TAPE divides the tasks into specific training, validation, and testing sets to ensure the evaluation of each task’s biological relevance and generalizability to real-life scenarios. TAPE serves as a benchmark for a range of semi-supervised protein representation learning methods, encompassing recent advances as well as established sequence-based learning techniques.
@@ 6. Genomic Large Language Models
iEnhancerBERT (Luo et al., 2022a)   iEnhancer-ELM (Li et al., 2023e)   miProBERT (Wang et al., 2023a)   TFBert (Luo et al., 2023a)   Enformer (Avsec et al., 2021)
@@
Nucleotide-Transformer (Dalla-Torre et al., 2023)   SA DNA-LM (Gankin et al., 2023)   GENA-LM (Fishman et al., 2023)   PLPMpro (Li et al., 2023b)   Uni-RNA (Wang et al., 2023b) ] [Other series: MoDNA (An et al., 2022)   ] ] [Decoder-only, text width=4.5em [LongConv series: HyenaDNA (Nguyen et al., 2023)   ] [GPT series: DNAGPT (Zhang et al., 2023e)   ] [Diffusion+Transformer series: GenSLMs (Zvyagin et al., 2022)   ] ] [Encoder-Decoder, text width=6em [ ENBED (Malusare et al., 2023)   ] ] ] [Datasets
@@
1000 Genomes Project (Consortium et al., 2015)   Panglao (Franzén et al., 2019)   ExPecto (Rusk, 2018)   BV-BRC (Olson et al., 2023)   RNAcmap (Rusk, 2018)   ] ] [Benchmark, text width=4em [CAGI5 (Katsonis and Lichtarge, 2019)   Protein-RNA (Horlacher et al., 2023)   NT-Bench (Dalla-Torre et al., 2023)   ] ] ] [Evaluation
@@ 6.1. Models
iEnhancer-ELM (Li et al., 2023e) 2022.12 110M BERT - Enhancer. pred. DNA
@@
SpliceBERT (Chen et al., 2023c) 2023.02 19.4M BERT UCSC Splicing pred. RNA
@@
DNAGPT (Zhang et al., 2023e) 2023.08 100M-3B GPT GRCh38 Func. pred. and generation DNA
@@
Encoder-only Models. In the realm of encoder-only architecture, significant models such as SpliceBERT (Zheng et al., 2019), DNABERT (Ji et al., 2021), DNABERT-2 (Zhou et al., 2023b), iEnhancer-BERT (Luo et al., 2022a), scBERT (Yang et al., 2022), BioSeq-BLM (Li et al., 2021), PLPMpro (Li et al., 2023b), GPN (Benegas et al., 2022), SA DNA-LM (Gankin et al., 2023), RNA-MSM (Zhang et al., 2023b), miProBERT (Wang et al., 2023a), TFBert (Luo et al., 2023a), MRM-BERT (Zhang et al., 2023a), GPN-MSA (Benegas et al., 2023a), and iDNA-ABF (Jin et al., 2022) are noteworthy attempts employing the Transformer encoder within genomics. These models utilize a mask training mechanism where partial gene sequences are masked, prompting the model to predict and complete them, progressively learning inherent patterns within gene sequences. They demonstrate versatility in fine-tuning for various gene sequence-related tasks like promoter prediction, transcription factor binding site prediction, and single nucleotide variability, achieving commendable performance. MoDNA (An et al., 2022) employs a Bert-like encoder-only architecture, introducing a unique training paradigm with a stacked Generator-Discriminator structure. During training, the Generator completes masked sections based on sequence and motif information, while the Discriminator assesses the completed sequences. This dual-layer architecture enables the simultaneous integration of sequence features and motif information, facilitating motif-oriented learning. GENA-LM (Fishman et al., 2023) introduces a suite of encoder-based foundational DNA language models designed for genomics research. These models, unique in their ability to handle long sequences of up to 36,000 base pairs, address the challenge of decoding genomic sequences, which requires understanding extensive contextual information across numerous nucleotides. The integration of a newly-developed recurrent memory mechanism allows these models to process even larger DNA segments. Nucleotide-Transformer (Dalla-Torre et al., 2023) is a foundation encoder-only model pre-trained on DNA sequences. It integrates data from 3,202 diverse human genomes and 850 genomes from various species. The study focuses on bridging the gap in predicting molecular phenotypes from DNA sequences, a task often limited by the scarcity of annotated data and difficulties in transferring learnings between different prediction tasks. These transformer models, with parameters ranging from 50M to 2.5B, provide transferable, context-specific representations of nucleotide sequences, enabling more accurate molecular phenotype predictions, even in scenarios with limited data. Remarkably, these models learn to focus on key genomic elements, such as enhancers that regulate gene expression, without supervision, and can help prioritize functional genetic variants. EpiGePT (Gao et al., 2023a) is a transformer-based language pretrained model designed for epigenomics. It predicts genome-wide epigenomic signals by considering transcriptional regulation mechanisms. EpiGePT uniquely incorporates context-specific activities of transcription factors, offering deeper insights into gene regulation compared to models trained only on DNA sequences. It shows exceptional performance in diverse epigenomic signal prediction tasks and can learn cell-type-specific long-range interactions through self-attention mechanisms. Uni-RNA (Wang et al., 2023b) is an encoder-only model that excels in predicting RNA structures and functions, outperforming previous methods in various tasks. It leverages a billion RNA sequences to extract hidden evolutionary and structural information, revolutionizing RNA research and facilitating drug development. As gene sequences often significantly exceed the length of natural language text, many models aim to overcome the quadratic time complexity of attention mechanisms. Models like Enformer (Avsec et al., 2021) and LOGO (Yang et al., 2021) employ a blend of convolutional downsampling to handle longer DNA sequences efficiently. Tools like BioSeq-BLM (Li et al., 2021) integrate traditional analysis methods with language models. PLPMpro (Li et al., 2023b) marks the advent of pre-training and fine-tuning in this domain, comparing language models’ effects with traditional methods.
@@
Decoder-only Models. Decoder-only models have gained significant attention due to their potent generative capacity. GenSLMs (Zvyagin et al., 2022) leverage genome-scale language models comprising multiple layers of attention-based decoders to elucidate the evolutionary dynamics of SARS-CoV-2. GenSLMs thus adeptly capture the evolutionary landscape of the SARS-CoV-2 genome. Pre-trained on prokaryotic gene sequences and further fine-tuned for SARS-CoV-2 genomes, GenSLM demonstrates precise and rapid identification of concerning variants. Besides commonly used architectures like transformers, niche architectures also find applications in genomic research. For instance, Models like DNAGPT (Zhang et al., 2023e) have introduced decoder-only architectures, akin to GPT, into the genomics domain by integrating specific symbols to guide different tasks, enabling zero-shot capabilities in species identification and regulatory factor prediction. HyenaDNA (Nguyen et al., 2023) incorporates a fast Fourier transform-accelerated recursive long-range convolution, a quadratic time complexity solution, enabling the handling of ultra-long DNA sequences. Employing convolution on the entire DNA sequence preserves single-nucleotide resolution, facilitating research into single nucleotide variability.
@@ 6.2. Datasets
Dataset Time Scale Keywords
@@
Currently, the majority of Gene-LLMs are trained on human DNA data. To increase the amount of training data and enhance the model’s generalization ability, genomes of organisms closely related to humans, such as chimpanzees or pigs, or even the genomes of all mammals are utilized for training. Here we list prevalent training datasets and benchmarks for Gene-LLMs, as summarized in Table 8.
@@ 6.3. Evaluation
  • Enhancer Prediction. Enhancers, short DNA regions, play a pivotal role in enhancing the likelihood of gene transcription by providing binding sites for various molecules, facilitating the assembly and stabilization of transcription machinery. Accurate prediction of enhancers holds significant importance in comprehending gene regulation, a critical aspect for understanding health and disease. Identifying enhancers can unveil genetic foundations of diseases and offer potential targets for therapeutic interventions. Notably, encoder-only models like iEnhancer-BERT (Luo et al., 2022a), iEnhancer-ELM (Li et al., 2023e) basically follow the same method, i.e., pretraining on human genome sequences and fintuning on enhancer data.
@@
Sequence Generation. Generation ability stands as a pivotal task in bioinformatics, focusing on models’ capacity to create artificial sequences that closely mimic real biological sequences. This ability becomes particularly crucial when considering the generation of artificial human genomes (AGs), serving as tools to safeguard genetic privacy and reduce costs linked with genetic sample collection. DNAGPT (Zhang et al., 2023e) emerges as a model purpose-built for this task. This model generated 5,000 AGs, each spanning a region of 10,000 Single Nucleotide Polymorphisms (SNPs). When compared with other models, DNAGPT showcased superior performance across various metrics.
@@ 7. Multi-modal Scientific Large Language Models
Multimodal Large Language Models have emerged as a prominent research area, utilizing robust LLMs as the core to handle multimodal data. These models possess a distinctive capability to incorporate diverse data types like text, images, audio, and other forms of information, enabling comprehensive exploration and problem-solving across various domains. The multimodal models hold promising prospects, particularly in the realms of biological and chemical sciences encompassing protein, molecular, and genomic studies. In this section, we explore recent advancements in multimodal models within these scientific fields (i.e., MM-Sci-LLMs), highlighting their capabilities and the exploited datasets. Note that, this survey focuses on cross-lingual multimodal models, involving at least two languages from different domains, such as text and molecule. Hence, we exclude the monolingual multimodal approaches from MM-Sci-LLMs, like the joint modeling of protein sequences and structures (Su et al., 2023; Wang et al., 2023d; Zhang et al., 2023c). The diverse forms of different languages and modalities are illustrated in Figure 4, and the overview of this section is shown in Figure 11.
@@
(Sec. 7.1),text width=4em [Molecule-Text, text width=5.5em [Encoder-only: Text2Mol (Edwards et al., 2021)   KV-PLM (Zeng et al., 2022)  MoMu (Zeng et al., 2022)  MolFM (Luo et al., 2023b) GPT-MolBERTa (Balaji et al., 2023) ] [Decoder-only: DrugChat (Liang et al., 2023)  MolReGPT (Li et al., 2023d)  MolXPT (Liu et al., 2023e) ] [Encoder-Decoder: MolT5 (Edwards et al., 2022)  Text+Chem T5 (Edwards et al., 2022)  ChatMol (Zeng et al., 2023b)  GIT-Mol (Liu et al., 2023b)
MoleculeSTM (Liu et al., 2023a) ] ] [Protein-Text, text width=5.5em [Encoder-only: ProTranslator (Xu and Wang, 2022)  ProtST-ProtBert (Xu et al., 2023b) ] [Decoder-only: InstructProtein (Wang et al., 2023c) ] [Encoder-Decoder: ProteinDT (Liu et al., 2023f)  Prot2Text (Abdine et al., 2023)  ProtST-ESM-1B (Xu et al., 2023b)  ProtST-ESM-2 (Xu et al., 2023b) ] ] [Protein-Molecule, text width=5.5em [Encoder-only: DrugCLIP (Gao et al., 2023b)] [Decoder-only: DrugGPT (Li et al., 2023a) ] [Encoder-Decoder: ChemBERTaLM (Uludoğan et al., 2022)  DeepTarget (Chen et al., 2023b) ] ] [Comprehensive, text width=5.5em [Encoder-only:BioTranslator (Xu et al., 2023a)] [Decoder-only: Galactica (Taylor et al., 2022b)  ChatDrug (Liu et al., 2023c)  DARWIN-MDP (Xie et al., 2023)
BioMedGPT-10B (Luo et al., 2023c) Mol-Instructions (Fang et al., 2023a) ] [Encoder-Decoder: BioT5 (Pei et al., 2023) ] ] ] [Datasets
(Sec. 7.2),text width=4em [Training, text width=3.5em [SwissProtCLAP (Liu et al., 2023f)   ProtDescribe (Xu et al., 2023b)  InstructProtein (Wang et al., 2023c)  Scientific Knowledge Dataset (Xie et al., 2023)
BioLip (Yang et al., 2012)  ChEBL-dia (Zeng et al., 2023b)  PubChemQA (Luo et al., 2023c)  Prot2Text (Abdine et al., 2023)  PCdes (Zeng et al., 2022)  PubChemSTM (Liu et al., 2023a)
MoMu (Zeng et al., 2022)  ChEBI-20 (Edwards et al., 2021)  Mol-Instructions (Fang et al., 2023a) ] ] [Benchmark, text width=3.5em [ MoleculeNet  (Wu et al., 2018)   BindingDB (Gilson et al., 2016)   DUD-E (Mysinger et al., 2012)  ] ] ] [Evaluation
@@
MolReGPT (Li et al., 2023d) 2023.06 - ChatGPT ChEBI-20
@@
MolXPT (Liu et al., 2023e) 2023.05 350M GPT-2 PubChem, PubMed $\times$
@@
GIT-Mol (Liu et al., 2023b) 2023.08 700M SciBERT, GIN, SwinTransformer (Liu et al., 2021a) PubChem, ChEBI-20 $\times$

### v2

@@ 5.2. Datasets
  • PIR (Wu et al., 2004).The Protein Information Resource(PIR) is a popular protein sequence database that provides information on functionally annotated protein sequences. PIR maintains three databases, the Protein Sequence Database (PSD), the Non-redundant Reference (NREF) sequence database, and the integrated Protein Classification (iProClass) database, which contains annotated protein sequences, classification information, and protein family, function, and structure information.
@@
  • HIPPIE (Schaefer et al., 2012). The HIPPIE database is a comprehensive database that encompasses information on protein-protein interactions in humans. The data originates from diverse sources, including experimental techniques such as yeast two-hybrid assays, mass spectrometry analysis, and documented interaction relationships reported in the scientific literature. These datasets are meticulously harmonized and standardized to provide precise and dependable insights into protein-protein interactions. This invaluable resource finds extensive applications in functional annotation, systems biology research, drug target discovery, and beyond.
@@
  • TAPE (Rao et al., 2019). The Tasks Assessing Protein Embeddings (TAPE) dataset is a collection of five biologically relevant semi-supervised learning tasks spanning different domains of protein biology, including secondary structure prediction, contact prediction, remote homology prediction, stability prediction, and fluorescence prediction. TAPE divides the tasks into specific training, validation, and testing sets to ensure the evaluation of each task’s biological relevance and generalizability to real-life scenarios. TAPE serves as a benchmark for a range of semi-supervised protein representation learning methods, encompassing recent advances as well as established sequence-based learning techniques.
  • Reactome (Jassal et al., 2020). Reactome is a database focused on human biomolecular reactions and pathways. It provides detailed networks of biochemical reactions, including proteins, small molecules, gene expression events, and their associated biological processes. Reactome can be used to validate the model’s predictions about the roles of proteins in biological processes.
  • STRING (Szklarczyk et al., 2023). STRING (Search Tool for the Retrieval of Interacting Genes/Proteins) is a comprehensive protein-protein interaction network database. It contains a wealth of information about known and predicted direct (physical) and indirect (functional) interactions. Through STRING, researchers can explore whether the proteins are in known interaction networks, or if they participate in unknown yet biologically plausible new interactions, which is crucial for understanding protein functions and complex biological pathways.
  • BioGRID (Oughtred et al., 2019). BioGRID (Biological General Repository for Interaction Datasets) is an online database that specializes in recording and curating protein-protein interactions, genetic interactions, chemical associations, and post-translational modifications from scientific literature. BioGRID serves as a critical tool for validating the predictive capabilities of these models regarding protein interactions and modifications. By leveraging the extensive interaction data in BioGRID, researchers can assess the accuracy of the model’s predictions about how proteins interact with each other and their roles in cellular processes.
  • InterPro (Paysan-Lafosse et al., 2023). InterPro is a comprehensive resource that offers functional analysis of protein sequences by classifying them into families and predicting the presence of domains and important sites. It integrates diverse information about protein families, domains, and functional sites from multiple reference databases. LLMs can utilize InterPro to annotate predicted protein sequences with potential functional domains, sites, and family memberships. Furthermore, InterPro’s comprehensive collection of protein signatures aids in validating and refining the predictions made by protein sequence models, ensuring a higher level of accuracy and reliability in protein function prediction and analysis.
  • DIP (Xenarios et al., 2002). The Database of Interacting Proteins(DIP) is a database that contains protein-protein interaction information that has been compiled through both manual curations and computational methods. It is useful for understanding protein functions, and their relationships with other proteins. It can also be used to study the properties of networks of interacting proteins, evaluate predictions of protein-protein interactions, and explore the evolution of these interactions.
  • PROSITE (Sigrist et al., 2012). The PROSITE is a collection of signatures that identify patterns or profiles in proteins, which can provide information on their biological functions. The signatures in the database are linked to annotation documents that provide information on the protein family or domain detected, including its name, function, 3D structure, and references.
@@ 6. Genomic Large Language Models
iEnhancerBERT (Luo et al., 2022a)   iEnhancer-ELM (Li et al., 2023g)   miProBERT (Wang et al., 2023a)   TFBert (Luo et al., 2023a)   Enformer (Avsec et al., 2021)
@@
Nucleotide-Transformer (Dalla-Torre et al., 2023)   SA DNA-LM (Gankin et al., 2023)   GENA-LM (Fishman et al., 2023)   PLPMpro (Li et al., 2023b)   Uni-RNA (Wang et al., 2023b)
FGBERT (Duan et al., 2024)   Geneformer (Cui et al., 2022)   RNAErnie (Wang et al., 2024)   gLM (Hwang et al., 2024)   RiNALMo (Penić et al., 2024)
GenomicLLM (Liu et al., 2024e) ] [SSM series: Caduceus (Schiff et al., 2024)   ] [Other series: MoDNA (An et al., 2022)   ] ] [Decoder-only, text width=4.5em [LongConv series: HyenaDNA (Nguyen et al., 2023)   Evo (Nguyen et al., 2024)   ] [GPT series: DNAGPT (Zhang et al., 2023f)   ] [Diffusion+Transformer series: GenSLMs (Zvyagin et al., 2022)   ] ] [Encoder-Decoder, text width=6em [ ENBED (Malusare et al., 2023)   ] ] ] [Datasets
@@
1000 Genomes Project (Consortium et al., 2015)   Panglao (Franzén et al., 2019)   ExPecto (Rusk, 2018)   BV-BRC (Olson et al., 2023)   RNAcmap (Rusk, 2018)   ] ] [Benchmark, text width=4em [CAGI5 (Katsonis and Lichtarge, 2019)   Protein-RNA (Horlacher et al., 2023)   NT-Bench (Dalla-Torre et al., 2023)   GenBench (Liu et al., 2024c)   BEACON (Ren et al., 2024)   ] ] ] [Evaluation
@@ 6.1. Models
iEnhancer-ELM (Li et al., 2023g) 2022.12 110M BERT - Enhancer. pred. DNA
@@
SpliceBERT (Chen et al., 2023e) 2023.02 19.4M BERT UCSC Splicing pred. RNA
@@
FGBERT (Benegas et al., 2023a) 2024.02 954M BERT - seq-func. relation DNA $\times$
gLM (Hwang et al., 2024) 2024.04 1B BERT - Func. pred. RNA
Caduceus (Schiff et al., 2024) 2024.04 1B SSM - Func. pred. RNA
RNAErnie (Wang et al., 2024) 2024.05 105M BERT - Func. pred. RNA
RiNALMo (Penić et al., 2024) 2024.05 650M BERT - Func. pred. RNA
@@
DNAGPT (Zhang et al., 2023f) 2023.08 100M-3B GPT GRCh38 Func. pred. and generation DNA
Evo (Nguyen et al., 2024) 2024.02 7B LongConv OpenGenome Pred. mol interactions. DNA $\times$
@@
Encoder-only Models. In the realm of encoder-only architecture, significant models such as SpliceBERT (Zheng et al., 2019), DNABERT (Ji et al., 2021), DNABERT-2 (Zhou et al., 2023b), iEnhancer-BERT (Luo et al., 2022a), scBERT (Yang et al., 2022b), BioSeq-BLM (Li et al., 2021), PLPMpro (Li et al., 2023b), GPN (Benegas et al., 2022), SA DNA-LM (Gankin et al., 2023), RNA-MSM (Zhang et al., 2023b), miProBERT (Wang et al., 2023a), TFBert (Luo et al., 2023a), MRM-BERT (Zhang et al., 2023a), GPN-MSA (Benegas et al., 2023a), iDNA-ABF (Jin et al., 2022) and FGBERT (Duan et al., 2024) are noteworthy attempts employing the Transformer encoder within genomics. These models utilize a mask training mechanism where partial gene sequences are masked, prompting the model to predict and complete them, progressively learning inherent patterns within gene sequences. They demonstrate versatility in fine-tuning for various gene sequence-related tasks like promoter prediction, transcription factor binding site prediction, and single nucleotide variability, achieving commendable performance. MoDNA (An et al., 2022) employs a Bert-like encoder-only architecture, introducing a unique training paradigm with a stacked Generator-Discriminator structure. During training, the Generator completes masked sections based on sequence and motif information, while the Discriminator assesses the completed sequences. This dual-layer architecture enables the simultaneous integration of sequence features and motif information, facilitating motif-oriented learning. GENA-LM (Fishman et al., 2023) introduces a suite of encoder-based foundational DNA language models designed for genomics research. These models, unique in their ability to handle long sequences of up to 36,000 base pairs, address the challenge of decoding genomic sequences, which requires understanding extensive contextual information across numerous nucleotides. The integration of a newly-developed recurrent memory mechanism allows these models to process even larger DNA segments. Nucleotide-Transformer (Dalla-Torre et al., 2023) is a foundation encoder-only model pre-trained on DNA sequences. It integrates data from 3,202 diverse human genomes and 850 genomes from various species. The study focuses on bridging the gap in predicting molecular phenotypes from DNA sequences, a task often limited by the scarcity of annotated data and difficulties in transferring learnings between different prediction tasks. These transformer models, with parameters ranging from 50M to 2.5B, provide transferable, context-specific representations of nucleotide sequences, enabling more accurate molecular phenotype predictions, even in scenarios with limited data. Remarkably, these models learn to focus on key genomic elements, such as enhancers that regulate gene expression, without supervision, and can help prioritize functional genetic variants. EpiGePT (Gao et al., 2023a) is a transformer-based language pretrained model designed for epigenomics. It predicts genome-wide epigenomic signals by considering transcriptional regulation mechanisms. EpiGePT uniquely incorporates context-specific activities of transcription factors, offering deeper insights into gene regulation compared to models trained only on DNA sequences. It shows exceptional performance in diverse epigenomic signal prediction tasks and can learn cell-type-specific long-range interactions through self-attention mechanisms. Uni-RNA (Wang et al., 2023b) is an encoder-only model that excels in predicting RNA structures and functions, outperforming previous methods in various tasks. It leverages a billion RNA sequences to extract hidden evolutionary and structural information, revolutionizing RNA research and facilitating drug development. Geneformer (Cui et al., 2022) is a context-aware, attention-based deep learning model developed for network biology applications with limited data. It is pretrained on a large-scale corpus of human single-cell transcriptomes, enabling it to make context-specific predictions. The model leverages transfer learning, allowing it to fine-tune and adapt to various downstream tasks, such as disease modeling and identification of therapeutic targets, with only a small amount of task-specific data. As gene sequences often significantly exceed the length of natural language text, many models aim to overcome the quadratic time complexity of attention mechanisms. Models like Enformer (Avsec et al., 2021) and LOGO (Yang et al., 2021) employ a blend of convolutional downsampling to handle longer DNA sequences efficiently. Tools like BioSeq-BLM (Li et al., 2021) integrate traditional analysis methods with language models. PLPMpro (Li et al., 2023b) marks the advent of pre-training and fine-tuning in this domain, comparing language models’ effects with traditional methods.
@@
Decoder-only Models. Decoder-only models have gained significant attention due to their potent generative capacity. GenSLMs (Zvyagin et al., 2022) leverage genome-scale language models comprising multiple layers of attention-based decoders to elucidate the evolutionary dynamics of SARS-CoV-2. GenSLMs thus adeptly capture the evolutionary landscape of the SARS-CoV-2 genome. Pre-trained on prokaryotic gene sequences and further fine-tuned for SARS-CoV-2 genomes, GenSLM demonstrates precise and rapid identification of concerning variants. Besides commonly used architectures like transformers, niche architectures also find applications in genomic research. For instance, Models like DNAGPT (Zhang et al., 2023f) have introduced decoder-only architectures, akin to GPT, into the genomics domain by integrating specific symbols to guide different tasks, enabling zero-shot capabilities in species identification and regulatory factor prediction. HyenaDNA (Nguyen et al., 2023) incorporates a fast Fourier transform-accelerated recursive long-range convolution, a quadratic time complexity solution, enabling the handling of ultra-long DNA sequences. Employing convolution on the entire DNA sequence preserves single-nucleotide resolution, facilitating research into single nucleotide variability. Evo (Nguyen et al., 2024) , a pretrained genome LLM with a modified StripedHyena architecture, is also capable of interpreting the core biological ”languages” of DNA, RNA, and proeins, allowing for both predictive analysis and generative design across molecular and genomic scales.
@@ 6.2. Datasets
Dataset Last updated Scale Keywords
@@
GenBench (Liu et al., 2024c) 2024.06 - Evaluating Genomics Foundational Models
BEACON (Ren et al., 2024) 2024.06 - Evaluating RNA Models
Currently, the majority of Gene-LLMs are trained on human DNA data. To increase the amount of training data and enhance the model’s generalization ability, genomes of organisms closely related to humans, such as chimpanzees or pigs, or even the genomes of all mammals are utilized for training. Here we list prevalent pre-training datasets and benchmarks for Gene-LLMs, as summarized in Table 8.
@@
  • GenBench: A Benchmarking Suite for Systematic Evaluation of Genomic Foundation Models (Liu et al., 2024c). GenBench is a benchmarking suite developed to provide a systematic evaluation framework for Genomic Foundation Models (GFMs). It addresses the challenge of equitable assessment by standardizing experimental settings, model complexities, benchmark datasets, and reproducibility. This suite includes evaluations across short- and long-range genomic tasks, covering key areas such as Coding Region, Non-Coding Region, and Genome Structure. GenBench supports a variety of state-of-the-art models, including DNABERT, HyenaDNA, and Nucleotide Transformer, among others. The benchmark spans 43 datasets, offering a thorough analysis of model performance and the interaction between model architectures and dataset characteristics. Through these evaluations, insights are gained into the strengths of different model types, guiding future GFM development.
  • BEACON: Benchmark for Comprehensive RNA Tasks and Language Models (Ren et al., 2024). This benchmark provides a thorough evaluation framework for RNA-related tasks, addressing the gap in standardized benchmarks for RNA models. BEACON assesses various RNA foundational models and traditional approaches, such as CNNs, ResNets, and LSTMs, across 13 tasks spanning structural analysis, functional studies, and engineering applications. The benchmark includes datasets from multiple sources like bpRNA, icSHAPE, and GENCODE, covering RNA sequence-level and nucleotide-level analyses. Key findings highlight the effectiveness of single nucleotide tokenization and the Attention with Linear Biases (ALiBi) positional encoding in RNA language models. BEACON also introduces a robust baseline model, BEACON-B, that demonstrates strong performance with limited data and computational resources.
@@ 6.3. Evaluation
  • Enhancer Prediction. Enhancers, short DNA regions, play a pivotal role in enhancing the likelihood of gene transcription by providing binding sites for various molecules, facilitating the assembly and stabilization of transcription machinery. Accurate prediction of enhancers holds significant importance in comprehending gene regulation, a critical aspect for understanding health and disease. Identifying enhancers can unveil genetic foundations of diseases and offer potential targets for therapeutic interventions. Notably, encoder-only models like iEnhancer-BERT (Luo et al., 2022a), iEnhancer-ELM (Li et al., 2023g) basically follow the same method, i.e., pretraining on human genome sequences and fintuning on enhancer data.
@@
Sequence Generation. Generation ability stands as a pivotal task in bioinformatics, focusing on models’ capacity to create artificial sequences that closely mimic real biological sequences. This ability becomes particularly crucial when considering the generation of artificial human genomes (AGs), serving as tools to safeguard genetic privacy and reduce costs linked with genetic sample collection. DNAGPT (Zhang et al., 2023f) emerges as a model purpose-built for this task. This model generated 5,000 AGs, each spanning a region of 10,000 Single Nucleotide Polymorphisms (SNPs). When compared with other models, DNAGPT showcased superior performance across various metrics.
@@ 7. Multi-modal Scientific Large Language Models
Multimodal Large Language Models have emerged as a prominent research area, utilizing robust LLMs as the core to handle multimodal data. These models possess a distinctive capability to incorporate diverse data types like text, images, audio, and other forms of information, enabling comprehensive exploration and problem-solving across various domains. The multimodal models hold promising prospects, particularly in the realms of biological and chemical sciences encompassing protein, molecular, and genomic studies. In this section, we explore recent advancements in multimodal models within these scientific fields (i.e., MM-Sci-LLMs), highlighting their capabilities and the exploited datasets. Note that, this survey focuses on cross-lingual multimodal models, involving at least two languages from different domains, such as text and molecule. Hence, we exclude the monolingual multimodal approaches from MM-Sci-LLMs, like the joint modeling of protein sequences and structures (Su et al., 2023; Wang et al., 2023h; Zhang et al., 2023d). The diverse forms of different languages and modalities are illustrated in Figure 4, and the overview of this section is shown in Figure 10.
@@
(Sec. 7.1),text width=4em [Molecule-Text, text width=5.5em [Encoder-only: Text2Mol (Edwards et al., 2021)   KV-PLM (Zeng et al., 2022)  MoMu (Zeng et al., 2022)  MolFM (Luo et al., 2023b) GPT-MolBERTa (Balaji et al., 2023) ] [Decoder-only: DrugChat (Liang et al., 2023)  MolReGPT (Li et al., 2023e)  MolXPT (Liu et al., 2023f)  DrugLLM (Liu et al., 2024b)  MolecularGPT (Liu et al., 2024a) ] [Encoder-Decoder: MolT5 (Edwards et al., 2022)  Text+Chem T5 (Edwards et al., 2022)  ChatMol (Zeng et al., 2023b)  GIT-Mol (Liu et al., 2023c)
MoleculeSTM (Liu et al., 2023b)  GIMLET (Zhao et al., 2023a)  Atomas (Zhang et al., 2024c)  3D-MolT5 (Pei et al., 2024b) ] ] [Protein-Text, text width=5.5em [Encoder-only: ProTranslator (Xu and Wang, 2022)  ProtST-ProtBert (Xu et al., 2023c) ] [Decoder-only: InstructProtein (Wang et al., 2023g)  ProLLaMA (Lv et al., 2024)  ProtLLM (Zhuo et al., 2024) ] [Encoder-Decoder: ProteinDT (Liu et al., 2023g)  Prot2Text (Abdine et al., 2023)  ProtST-ESM-1B (Xu et al., 2023c)  ProtST-ESM-2 (Xu et al., 2023c)
PAAG (Yuan et al., 2024)  ProtT3 (Liu et al., 2024d) ] ] [Protein-Molecule, text width=5.5em [Encoder-only: DrugCLIP (Gao et al., 2023b)] [Decoder-only: DrugGPT (Li et al., 2023a) ] [Encoder-Decoder: ChemBERTaLM (Uludoğan et al., 2022)  DeepTarget (Chen et al., 2023c) ] ] [Gene/Cell-Text, text width=5.5em [Decoder-only: Cell2Sentence (Levine et al., 2023) GenePT (Chen and Zou, 2023) ChatCell (Pei et al., 2024c) GPTCelltype (Hou and Ji, 2024)] [Encoder-Decoder: CellPLM (Chang and Ye, 2024) CellWhisperer (Schaefer et al., [n. d.]) LangCell (Zhao et al., 2024b) ] ] [Comprehensive, text width=5.5em [Encoder-only:BioTranslator (Xu et al., 2023b)  BioBridge (Wang et al., 2023e)] [Decoder-only: Galactica (Taylor et al., 2022b)  ChatDrug (Liu et al., 2023d)  DARWIN-MDP (Xie et al., 2023)
BioMedGPT-10B (Luo et al., 2023c)  Mol-Instructions (Fang et al., 2023a)  ChemDFM (Zhao et al., 2024a)  ChemLLM (Zhang et al., 2024b) ] [Encoder-Decoder: BioT5 (Pei et al., 2023)  BioT5+ (Pei et al., 2024a) ] ] ] [Datasets
(Sec. 7.2),text width=4em [Training, text width=3.5em [SwissProtCLAP (Liu et al., 2023g)   ProtDescribe (Xu et al., 2023c)  InstructProtein (Wang et al., 2023g)  Scientific Knowledge Dataset (Xie et al., 2023)
BioLip (Yang et al., 2012)  ChEBL-dia (Zeng et al., 2023b)  PubChemQA (Luo et al., 2023c)  Prot2Text (Abdine et al., 2023)  ProteinLMDataset (Shen et al., 2024)  PCdes (Zeng et al., 2022)  PubChemSTM (Liu et al., 2023b)
MoMu (Zeng et al., 2022)  ChEBI-20 (Edwards et al., 2021)  Mol-Instructions (Fang et al., 2023a)  GEO (Clough and Barrett, 2016)  HCA (Regev et al., 2017)  ARCHS4 (Lachmann et al., 2018)
NCBI (Wheeler et al., 2007)  cellxgene (Megill et al., 2021)  SRT (He et al., 2021a)  CellTypist (Domínguez Conde et al., 2022)  scLibrary (Zhao et al., 2024b)  ] ] [Benchmark, text width=3.5em [ MoleculeNet  (Wu et al., 2018)   BindingDB (Gilson et al., 2016)   DUD-E (Mysinger et al., 2012)  ProteinLMBench (Shen et al., 2024) ] ] ] [Evaluation
@@
GIMLET (Zhao et al., 2023a) 2023.06 - T5 Chembl
MolReGPT (Li et al., 2023e) 2023.06 - ChatGPT ChEBI-20
@@
MolXPT (Liu et al., 2023f) 2023.05 350M GPT-2 PubChem, PubMed $\times$
@@
GIT-Mol (Liu et al., 2023c) 2023.08 700M SciBERT, GIN, SwinTransformer (Liu et al., 2021a) PubChem, ChEBI-20 $\times$